40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 23 23:10:19 2014, ericq , Update, General, Vent so far 
    Reply  Mon Jan 27 21:36:35 2014, ericq , Update, General, Vent so far 
       Reply  Tue Jan 28 22:24:46 2014, ericq , Update, General, Further Alignment 
Message ID: 9572     Entry time: Thu Jan 23 23:10:19 2014     Reply to this: 9579
Author: ericq  
Type: Update 
Category: General 
Subject: Vent so far 

[ericq, Manasa, Jenne]

Summary: We opened up the BS and both ITM chambers today, and put the light doors on. //Edit : Manasa  Post-vent the MC was very much misaligned in yaw. Both the ITMs moved in pitch as inferred from the oplev; but there is still light on the oplev PDs//. We toiled with the PMC and mode cleaner for a while to get reasonable transmission and stability (at least for a period of time). We then tried to lock IR to the y-arm, to no avail. 

Locking the PMC doesn't seem very robust with the low power level we have; adjusting the gain at all when it's locked throws it right out. The mode cleaner spot was visibly moving around on MC2 as well. We'll continue tomorrow. 

Details about alignment efforts: Manasa and I tried for a while to try and align the y-arm for IR. Straight out of venting the green TM00 would lock to the y-arm with about .45, as compared to .8 before venting, so it didn't seem to drift too far. The x-arm would even flash any modes, however. For a while, IR was no where to be seen after the mode cleaner. Eventually, we used the tip tilts to bring the AS beam onto the camera, which exhibited fringes, so we knew we were hitting the ITMs somewhere. We wandered around with the ETM to see if any retroflection was happening, and saw the IR beam scatter off of the earthquake stop. We moved it to the side to see it hitting the OSEM holder, and moved down to the bottom OSEM holder to get an idea of where to put pitch to get roughly the center of the ITM, then undid the yaw motion.

There, we would see very infrequent, weak flashes. We weren't able to distinguish the mode shape though; however, the flashes were coincident with where the green would lock to a very yaw-misaligned fishbone mode, to the lower right of the optic's center. We figured that if we gradually fixed the green alignment with the mode shapes we could see and actually lock on, we could use the tip tilts to adjust the IR pointing and keep it coincident and eventually resonate more. However, this didn't really work out. The flashes were very infrequent, and at this point the PMC/MC were getting very touchy, and would cease to stay locked for more than a minute or two. At this point, we stopped for the day. 

 

ELOG V3.1.3-